OpenStudy (anonymous):

If f(x) = x^4 + 4, g(x) = x - 7 and h(x) = \sqrt(x) , then f(g(h(x))) =

OpenStudy (praxer):

$$h(x)= \sqrt{x}$$ $$g(x)=x-7$$ $$f(x)=x^4+4$$ here for $$g(h(x))=g(\sqrt{x})=\sqrt{x}-7$$ again $$f(g(h(x)))=f(g(\sqrt{x})=f(\sqrt{x}-7)$$ so, the solution is $$f(\sqrt{x}-7)=(\sqrt{x}-7)^4+4$$